3x^2+16=-14x in AC method
Answer by Boreal(15235)

3x^2+14x+16=0
(3x+ )(x + )
write as
3x^2+6x+8x+16=0
3x(x+2)+8(x+2)=0
(3x+8((x+2)=0
x=-8/3, -2